OpenStudy (anonymous):

what are the rectangular coordinates of (-4, 7pi/6)

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} (&-4 &\frac{7\pi}{6})\\ &\uparrow &\uparrow \\ &r&\theta \end{array}\qquad \begin{array}{ccllll} x=rcos(\theta)\qquad &y=rsin(\theta)\\ \quad \\ x=(-4)cos(\frac{7\pi}{6})&y=(-4)sin(\frac{7\pi}{6}) \end{array}\)